OpenStudy (anonymous):

prove n! >= 5^n for all n >= 12

OpenStudy (freckles):

so you proved it for the base case?

OpenStudy (freckles):

that is you proved it for n=12?

OpenStudy (freckles):

anyways after that suppose k!>=5^k for some integer k>=12 and show (k+1)!>=5^(k+1) also holds.

OpenStudy (anonymous):

My bad, yeah 12 is the base case

OpenStudy (freckles):

so you have no problem with that part right?

OpenStudy (anonymous):

I get the base case yeah, and I understand I need to show (k+1)!>=5^(k+1) but I have no idea from there

OpenStudy (freckles):

Ok so our k's our only from 12 to bigger so for any value k we have that k>5 right?

OpenStudy (freckles):

Take the inequality k!>=5^k and see what happens if you multiply both sides by 5

OpenStudy (anonymous):

Yes but then once you get to that part (that is 5(n!) >= 5(5^n) don't you then have to show that you're after (n+1)! >= 5(5^n)

OpenStudy (freckles):

5(5^n) is 5^(n+1) now we just said k>5 if k>5 then k+1>5

OpenStudy (freckles):

so that means (k+1)*k!>=5*k! correct?

OpenStudy (anonymous):

Yeah okay, and why is k>5 for any value k?

OpenStudy (freckles):

recall that k is bigger than or equal to 12

OpenStudy (anonymous):

Oh I see where you're coming from

OpenStudy (anonymous):

so k will always be more than 5

OpenStudy (freckles):

yep

OpenStudy (freckles):

since k>=12

OpenStudy (anonymous):

So does (k+1)*k!>=5*k! prove it?

OpenStudy (freckles):

and write what we had on the other side of the 5*k!

OpenStudy (freckles):

\[(k+1)\cdot k! \ge 5 k! \ge 5(5^k)\]

OpenStudy (anonymous):

I think I see now, you've been a great help! Thanks!

OpenStudy (freckles):

np